Calculate the molarity, molality and mass percent of an NH3 solution made up of 30.0 g of NH3 in 70.0 g of water. The density of the solution is 0.982 g/mL


1
Expert's answer
2021-02-18T07:51:46-0500

No. Of moles of NH3 = 30/17 = 1.764 moles

Mass of water = 70 g

Mass of solution = 100g

Volume of solution = 100×0.982 = 98.2 ml

Volume of solvent = 70ml

Molarity = (1.764/98.2)×1000 = 17.96

Molality = (1.764/70)×1000 = 25.2

Mass % NH3 = (30/100)×100 = 30%

Mass % of water = (70/100)×100 = 70%